OpenStudy (anonymous):

What are the x- and y-intercepts of the graph of y = 2x^2 - 8x - 10 x-intercepts: (1, 0) and (-5, 0) y-intercept: (0, 10) x-intercepts: (-1, 0) and (-5, 0) y-intercept: (0, 5) x-intercepts: (-1, 0) and (5, 0) y-intercept: (0, 5) x-intercepts: (-1, 0) and (5, 0) y-intercept: (0, -10)

myininaya (myininaya):

To find x-intercepts, set y=0 and solve for x To find y-intercepts, set x=0 and solve for y

OpenStudy (anonymous):

meaning without doing much work at all you can rule out the top three

myininaya (myininaya):

because finding y-intercept is easiest! :)

myininaya (myininaya):

set x=0, what does y=?

OpenStudy (ash2326):

y=2x^2-8x-10 for y intercept put x=0 y=-10 so y intercept is (0,-10) for x intercept put y=0 and solve for x 0=2x^2-8x-10 divide by 2 the whole equation x^2-4x-5=0 find factors of -5 which have a difference of -4 -5 and 1 x^2-5x+x-5=0 x(x-5)+1(x-5)=0 ( x+1) (x-5)=0 x=-1 and 5 so x intercept (-1,0) and (5,0)